Reports until 17:58, Tuesday 17 December 2024
GQuEST Output Filter Cavities
Jeffrey Wack - posted 17:58, Tuesday 17 December 2024 (12055)
OFC1 Cavity Realignment (continued)

[Torrey, Jeff]

Today we continued the re-alignment of output filter cavity one. The 775 path is now aligned well enough to lock, and the 1550 path is aligned well enough to see flashes in transmission but still needs input alignment tuning to suppress the higher-order modes.

As discussed in [12053], we decided to realign the first output filter cavity in order to avoid the 775 beam clipping the piezo mount on transmission. It was clear that the piezo controlled mirror was a bit crooked (pitched down), and this required a wonky input alignment which led to clipping. We dissasembled and reassembled the piezo optic assembly and found that we could change the tilt of the mirror significantly by changing the compression of the viton by a very small amount, from which we conclude that the viton is being scrunched and we had perhaps been over-tightening it. We also experimented with a few rotations of the assembly, and the configuration we settled on was a 90 degree rotation from where we started. After this the mirror was noticably less wonky, and we proceeded with realigning the cavity, doing the 775 path first as it is more geometrically constrained by exiting through the piezo.

We performed a coarse alignement by removing the 775 input mirror and then using the cavity mirrors to postion the round-trip beam to be coincident with the incoming beam at approximately the location of the input mirror. We then used an iris to mark the round trip beam after it leaves the cavity, and after replacing the input mirror we tuned it such that the prompt reflection hit the same iris. This worked well enough to see a second pass beam in transmission, from which we could do further alignment.

We tuned up the cavity and input alignment to maximize power in the 775 0,0 mode (see attatched scan on the photodiode).

Next we roughly aligned the 1550 in an analogus way by removing the input mirror, etc. We are now seeing lots of modes flashing during a scan, and will continue to tune up the input alignment to eliminate the higher order modes. We also started setup for a fiber coupler to send the light to the second filter cavity, using a flipper mirror to toggle between diagnostic (camera and PD) and fiber coupler. (see attatched diagram)

Images attached to this report
Non-image files attached to this report